OpenStudy (anonymous):

how would I simplfy this (cos x)+(sin x+2)-(sinx-4)(cos x) which should equal 6cosx?

OpenStudy (anonymous):

oops it actually (cos x)(sin x+2)-(sinx-4)(cos x) which should equal 6cosx

OpenStudy (phi):

if the expression were x*(y+2) - (y-4)*x could you simplify it?

OpenStudy (anonymous):

yes! However my problem is I don't know the tricks yet on simplifying trig function

OpenStudy (phi):

No trick with this one. treat cos(x) as one thing (just like x in my problem). same with sin(x) I'll do the first term: cos(x)*( sin(x)+2) = cos(x)*sin(x) + 2*cos(x) Just like x(y+2)= xy +2x

OpenStudy (anonymous):

Sin (x)-4*Cos(x)= Cos(x)*Sin(X)-4*Cos(x)

OpenStudy (phi):

2 things: you should keep the parentheses (sin(x) - 4)*cos(x) because this is very different from sin(x)-4*cos(x) Also, the 2nd term has a minus sign out front. Don't forget it.

OpenStudy (anonymous):

oh, so the cosxsinx cancel out and we are left with 2cosx-(-4cosx)

OpenStudy (phi):

yes. and - - 4cos(x) becomes +4cos(x)

OpenStudy (anonymous):

which 6cosx

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thanks!